A random sample of=49 observations is drawn from a population with a mean equal to 20

and a standard deviation equal to14.

a. Find the probability that x overbarx is less than 14.

b. Find the probability that x overbarx is greater than 25.

c. Find the probability that x overbarx falls between 14 and 22

a. The probability that x overbarx is less than 14 is

_________​(Round to three decimal places as​ needed.)
Answer #1

SOLUTION:-

Given data,

n=49 ,m= 20, s=14.

a. Find the probability that x overbarx is less than 14:-

P(x<14)=P(Z<rac{x-mu }{s})

-б 14

P(z < 14) P(Z <-0.43)

From Z table

P(Z <-0.43) = 0.3336

b. Find the probability that x overbarx is greater than 25:-

P(X > 25)-1- P(Z< 25)

2520 14 P(X > 25) = 1-P(Z <

14

P(X > 25) 1- PZ 0.36)

P(X > 25) 1-0.6406

P(X > 25) 0.3594

c. Find the probability that x overbarx falls between 14 and 22:-

P(14<X<22)=P(z<22)-P(z<14)

P(14<X<22)=P(z<rac{22-20}{14})-P(z<rac{14-20}{14})

P(14<X<22)=P(z<(rac{2}{14})-P(z<rac{-6}{14})

P(14<X<22)=P(z<0.14)-P(z<-0.43)

P(14 < X < 22) = 0.4447-0.3336

P(14<X<22)=0.1111​​​​​​​